The degree of dissociation of water at 25°C is 1.9*10^-7 % and density is 1g-cm^-3. The ionic constant for water is?

Grade:11

1 Answers

Elanji Rahul
6 years ago
Ionic constant (or)product=xalpha=1.9×10^-7H20~h+ohH+=c×alphaMolarity=weight/gram molecular weight ×volumeWeight=1000th (1000ml)=1000/18=55.5 MOLARalpha=h+/c1.9×10^-7=x.18, 10^3Therefore, x=10^-7